OpenStudy (anonymous):

write the expression as the sine, cosine or tangent of a single angle cos8x cos6x - sin8x sin6x

random231 (random231):

this comes under the direct tansformation of cos(a+b)

OpenStudy (mathmale):

Does the following remind you of anything (a trig identity?) ??\[\cos a \cos b - \sin a \sin b\]

random231 (random231):

are you just here to get the answer? if that is so OS isnt the right place for you!

OpenStudy (dumbcow):

everything you need is right here http://en.wikipedia.org/wiki/Trig_identity#Angle_sum_and_difference_identities